OpenStudy (anonymous):

how many cm are in 25km?

OpenStudy (anonymous):

2500000 cm

OpenStudy (frostbite):

Remmeber that km is shorten for kilometer. kilo meaning 1000. So one km is the same as \(10^{3}m\). From m to cm we get \(1 m = 10^{2} cm\) So the calculator is going to be: \[25 km \times 10^{5} \frac{ cm }{ km }= 25 \times 10^{5} cm=2500000 cm\]

OpenStudy (frostbite):

Calculation*